# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

### v2

@@ Abstract
Nets, cooperative networks of neurons, have been proposed as format for the representation of sensory signals, as physical implementation of the Gestalt phenomenon and as solution to the neural binding problem, while the direct interaction between nets by structure-sensitive matching has been proposed as basis for object-global operations such as object detection. The nets are flexibly composed of overlapping net fragments, which are learned from statistical regularities of sensory input. We here present the cooperative network architecture (CNA), a concrete model that learns such net structure to represent input patterns and deals robustly with noise, deformation, and out-of-distribution data, thus laying the groundwork for a novel neural architecture.
Keywords net fragments  $\cdot$ neural code  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neural networks  $\cdot$ cooperativity  $\cdot$ neuroscience  $\cdot$ machine learning
@@ 1 Introduction
Figure 1: Representation of Objects by Nets. Objects are captured in their entirety by coherent nets that are composed of overlapping net fragments (schematically symbolized by white circles in the right panel). After training on natural images, each patch of visual space contains a complement of net fragments that represent textures that have been encountered with statistical significance. Net fragments overlap neuron-wise, and those activated by an object coalesce into a coherent net. The conundrum (stated in Olshausen and Field (2005)) that object contours can often not be found with the help of edge detectors due to lack of gray-level contrast (as in places inside the square in the left panel) may find its solution in that contours are defined as borders of the nets covering objects (or covering the background). Figure adapted from Olshausen and Field (2005), with permission.
Artificial intelligence (AI) systems have demonstrated remarkable capabilities in processing diverse types of data, including images, text, and speech (LeCun et al., 2015; Simmler et al., 2021). While many recent AI systems predominantly favor neural representations over symbolic ones, we follow a proposal by von der Malsburg et al. (2022) and present the cooperative network architecture (CNA), an architecture that uses structured networks of mutually (‘recurrently’) connected neurons as representation of sensory signals. Using structured networks to encode information brings a diverse set of advantages, including high robustness, ability of figure reconstruction, and representing of out-of-domain data, laying the foundations to build novel object recognition frameworks (see Section 6).
The networks in the CNA are called nets and are dynamically assembled from overlapping net fragments, where net fragments are connectivity patterns that are learned by synaptic plasticity based on the statistics of sensory input signals. A major argument in favor of this data architecture is the possibility (as described repeatedly by Hinton (1981); von der Malsburg (1981); Anderson and Van Essen (1987); Kree and Zippelius (1988); Olshausen et al. (1995); Wiskott et al. (1997); Wolfrum et al. (2008); Memisevic and Hinton (2010); Arathorn et al. (2013)) of solving the invariant object recognition problem by structure-preserving mapping between nets representing variant images and invariant models of objects. The representation of image structure by recurrently connected nets formed between elementary feature detector neurons is a fundamental deviation from the current artificial neural networks (ANNs) paradigm, according to which image structure is represented by sets of complex-feature neurons activated by purely feedforward connections (see Section 2.3 for a comparison of the two approaches).
In the cited models of invariant object recognition, the formation of object-representing networks within the image domain (the primary cortical area where sensory signals arrive – here called ‘area’) has been sidestepped by letting spatial neighborhood stand in for connectedness. Neighborhood, however, is too weak a condition to qualify neurons as having responded to the same object (which is, for instance, apparent when looking at an object through foliage). Following von der Malsburg et al. (2022), we rather assume that for sets of neurons to be admitted into a net, they have to have been co-activated frequently in the past by structured visual input and consequently have been tied together by Hebbian synaptic plasticity (Hebb, 1949) into net fragments. Net fragments play a role somewhat related to the code-book vectors of some image compression schemes (Taubman and Marcellin, 2002) in representing statistically dominant texture patterns within any patch of the visual field; the formation of net fragments is somewhat related to, though fundamentally different from, associative memory (Hopfield, 1982), see Section 2.1.
CNA maintains that when a region of the area is hit by signals caused by an object, it transiently activates a subset of all the neurons in the area. The majority of those neurons are not part of a co-activated net fragment and have to be silenced. This is achieved with the help of some periodically rising inhibition, which silences those neurons and leaves active the small subset of all transiently activated neurons that are supporting each other by excitatory connections. Due to neuron-wise overlap, net fragments merge into a coherent net, which can span the entire region covered by an object’s image and thus connect feature neurons that are spatially distant through chains of locally connected intermediary neurons, see Figure 1. It is this connectedness that unites neurons activated by an object in forming a structured net that separates those neurons from the background and from occluding elements of other objects, thus solving the segmentation problem (von der Malsburg, 2024) and tying them into a network amenable to be mapped to an invariant model network in another cortical area (see the outlook in Section 6).
This study introduces the cooperative network architecture (CNA), representing the first concrete implementation of some fundamental components proposed in von der Malsburg et al. (2022). We focus on how to leverage networks as neural code and, specifically, discuss the
  • establishment of a computational framework comparable to recent machine learning advances,
  • learning of net fragments,
  • demonstration of the activation of object-spanning networks (nets),
  • demonstration of filter properties afforded by nets,
  • discussion of nets in contrast to established ideas in neural computation and machine learning,
while relating to the significance of nets for invariant object recognition in multi-area settings only as outlook in the discussion section (i.e. we focus here on a single area while object recognition based on multiple areas is proposed for future research). We demonstrate that net fragments can be learned without relying on a teaching signal, whether supervised or unsupervised. This neural code exhibits robust capabilities for handling out-of-distribution data and effectively represents objects composed of fundamental features encountered during training—a task at which autoencoders (Hinton and Salakhutdinov, 2006) fail.
2 Relation to Alternative Concepts
The model we describe below is the first to demonstrate in concrete form the generation of structured nets that form a neural code to represent visual structure. To grasp the significance of the approach, it is important to compare it to alternative approaches.
2.1 Associative Memory
Associative memory (AM), as formulated by Amari (1972) or Hopfield (1982), shares with the present study the aim of retrieving previously encountered structure, and in both, the structure is recorded with the help of Hebbian synaptic plasticity. Whereas AM deals with the storage and retrieval of neural firing patterns, our purpose is to activate networks, that is, connectivity patterns that are structured by previously recorded connectivity pattern fragments. Admittedly, in both cases, the (re)construction process results in the activation of subsets of all neurons of the area, but whereas in AM, the connectivity of the set of active neurons is all-to-all, in our case, connectivity is restricted to within the net fragments that have previously been recorded, which in our visual application are restricted to short-range (local) connections and express two-dimensional topological structure. It is this structured connectivity that serves as basis for the structure-sensitive matching mentioned in the introduction. The monolithic all-to-all connectivity of the AM, by distinction, does not afford such structure-sensitive matching (i.e. instead of representing each structure as monolithic whole, the CNA learns to decompose them into local sub-patterns (net fragments), which can then be dynamically composed into nets to represent structures).
This functional difference is conditioned on a different mode of learning. Whereas in AM, a neural pattern is stored in a single event by strengthening the connections between any pair of currently active neurons, our approach distinguishes significant from insignificant connections, significant being those that are strengthened again and again, while insignificant are those that are strengthened only occasionally. This learning is driven by exposure to a large number of patterns, none of which may ever repeat, but which contain small repeating subpatterns. It is these that generate net fragments. In the visual case, it is local texture patches that play this role of repeating sub-patterns.
A basic limitation of AM is that unless the patterns to be stored are statistically independent (orthogonal), that is, have small neuron-wise overlap, they interfere with each other by the spurious activation of neurons. The problem can be reduced (as proposed by Tsodyks and Feigel’man (2007) or Palm (2013)) if patterns are sparse, that is, activate only a small fraction of all neurons. The statistical mode of pattern storage proposed here opens a more fundamental solution to the problem of pattern overlap. We propose that each input feature (specified by feature type and position) is represented not by one neuron but by a number of ‘alternative neurons’, neurons that have the exact same receptive field (for examples, see Cossell et al. (2015)), and thus are activated by the same afferents  [^1] There is ample space for alternative neurons, as there are 100 times more neurons in primary visual cortex than output neurons in the retina (Leuba and Kraftsik, 1994).. While these neurons have the same afferent input, they can exhibit different excitatory output. Given this expansion, different net fragments with feature-wise overlap can reduce neuron-wise overlap by making use of different alternative neurons.
AM and our approach differ in yet another aspect, the mode of retrieval (or, in our case, construction) of patterns. In the version of Hopfield (1982) of AM, this is realized by minimizing an energy function by sequential switching of neurons. In our approach, winning patterns are projected out from the set of all neurons activated by the input and by rising inhibition that silences all those alternative neurons that do not have sufficient recurrent excitatory support, no neuron-sequential process being necessary.
2.2 Boltzmann Machine
Some important aspects of the system illustrated by our model are highlighted best by comparing it with the Boltzmann Machine (BM) proposed in Hinton and Sejnowski (1983). Both systems share features such as the ability to admitting recurrent connections (as does Mumford (2003)), but they also exhibit significant differences. Nevertheless, as our focus is limited to one area and is not the full architecture as proposed in von der Malsburg et al. (2022), we highlight only select differences here.
The BM makes a clear distinction between visible and hidden (‘latent’) neurons. In the CNA this distinction is softened: it posits neurons that are activated by sensory input, but of which the majority are quickly silenced by inhibition in favor of those neurons that support each other through previously learned connections. A winning alternative neuron can be interpreted as the hypothesis not only that the neuron has detected its trigger pattern in its afferents but also that the neuron has connections to a number of other active neurons and is consistent with them in light of previous experience.
In both BM and our model, the activity of hidden, or in our case alternative, neurons, is controlled by the same type of dynamic equations, with the difference that we do not insist on symmetric connections and therefore do not derive an energy function and consequently cannot prove convergence of the neural dynamics (see, however, the remark made in Hinton and Sejnowski (1983), that given sufficiently redundant support of active neurons, the missing of up to half of the connections due to lack of symmetry does not jeopardize stability of a state). In a single cycle of the periodically rising inhibition, convergence is not an issue, as alternative neurons undergo winner-take-all competition.
2.3 Artificial Neural Networks (ANN)
In the concept pursued here, objects and other structures within the input are represented by nets, sets of neurons that are held together by cooperative excitatory connections and are activated as a whole, the activation of individual neurons being conditioned on the activation of other neurons in the same net. In contrast, a large variety of ANN versions and learning techniques have been proposed, including multilayer perceptrons (MLPs) (Prince, 2023), transformers (Vaswani et al., 2017), autoencoders (Hinton and Salakhutdinov, 2006), or representation learning (Bengio et al., 2014). Common to all of these is that the sensory input is converted through many layers into a representation in the form of a set of neuron states. These neurons can be considered as propositions that indicate the presence of complex features (activity patterns) within the input. This representation of the input is then exploited by evaluation circuits (often in the form of a few fully connected layers of neurons) to achieve functions such as object classification, next-token prediction, or, in autoencoders, reconstruction of the input. The common feature of all ANN systems is that the connections from the input to the representation layer and from there to the output are purely feedforward and are trained by back-propagation of error (Rosenblatt, 1962, p. 292).
The CNA proposes a clear separation between input representation (within an area) and its functional evaluation (between areas, see Section 6). The representation of the input is trained by its statistics and is thus essentially neutral and without specialization to any specific function, giving it foundational character. In ANN systems, the selection of patterns to be extracted from the input is strongly biased to serve the trained function, which may be the reason that a transformer trained on one set of functions cannot continue to be trained on another (Dohare et al., 2024).
ANN’s hierarchy of layers of neurons responding to features of intermediate complexity is replaced in the CNA by a single layer, in which net fragments of different sizes represent a hierarchy of features of varying complexity (and inter-mesh to form coherent nets standing for whole objects, thus solving the binding problem). The purpose of transition to other areas is, in the CNA, not transition to higher complexity but transition to representations that generalize over, e.g., position, size, or orientation of nets in lower areas, as discussed in Section 6.
In ANN systems, the high-level feature neurons that are activated by an input do not express any relations between them. This ‘binding problem’ (von der Malsburg, 1981; Burwick, 2014; Greff et al., 2020)) can give rise to ambiguity, e.g., the hallucination of feature conjunctions (Treisman and Schmidt, 1982). This ambiguity can be reduced by additional neurons whose trigger patterns overlap those of other neurons whose relative position would otherwise be in doubt (autoencoders (Hinton and Salakhutdinov, 2006) show that ambiguity can be suppressed altogether by a sufficiently large number of hidden neurons). In the CNA, the binding problem is solved in a natural way by net fragments that interconnect sets of features that co-occur with statistical significance in the input and that merge together into nets representing coherent chunks within a given input. This simplicity of representation of visual structure in CNA is due to its fundamentally different perspective on neural representation, in that the active state is not only described by the elementary features that are active but also by the connections between them. This is made possible by different alternative neurons that represent the same elementary feature but express different connectivity.
3 Computational Framework

### v3

@@ Abstract
We introduce the Cooperative Network Architecture (CNA), a model that represents sensory signals using structured, recurrently connected networks of neurons, termed “nets.” Nets are dynamically assembled from overlapping net fragments, which are learned based on statistical regularities in sensory input. This architecture offers robustness to noise, deformation, and out-of-distribution data, addressing challenges in current vision systems from a novel perspective. We demonstrate that net fragments can be learned without supervision and flexibly recombined to encode novel patterns, enabling figure completion and resilience to noise. Our findings establish CNA as a promising paradigm for developing neural representations that integrate local feature processing with global structure formation, providing a foundation for future research on invariant object recognition.
Keywords net fragments  $\cdot$ nets  $\cdot$ neural code  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neural networks  $\cdot$ machine learning
@@ 1 Introduction
Figure 1: Representation of Objects by Nets. Objects are captured in their entirety by coherent nets composed of overlapping net fragments (schematically symbolized by white circles in the right panel). After training on natural images, each patch of visual space contains a complement of net fragments that represent textures that have been encountered with statistical significance. Net fragments overlap neuron-wise, and those activated by an object coalesce into a coherent net. The conundrum (stated in Olshausen and Field (2005)) that object contours can often not be found with the help of edge detectors due to lack of gray-level contrast (as in places inside the square in the left panel) may be resolved by the idea that contours are defined as borders of the nets covering objects (or covering the background). Figure adapted from Olshausen and Field (2005), with permission.
Artificial intelligence (AI) systems have demonstrated remarkable capabilities in processing diverse types of data, including images, text, and speech (LeCun et al., 2015; Sager et al., 2025). Yet, despite these advances, artificial vision systems continue to struggle with robust object representation and recognition, particularly under varying viewing conditions, occlusion, or noise (Qi et al., 2023; Simmler et al., 2021). This limitation highlights a fundamental gap between current AI systems and the flexible, context-aware perception observed in biological vision.
Existing approaches attempt to address these challenges through methods such as feature or sample engineering (Amirian et al., 2018; Moosavi-Dezfooli et al., 2016), domain adaptation (Csurka, 2017; Sager et al., 2022), and data augmentation (Carlucci et al., 2019; Tuggener et al., 2024). However, these techniques largely operate within the conventional paradigm of representing objects as static patterns of neuronal activations, often requiring large datasets and extensive supervision to achieve generalization.
In contrast, we propose a new paradigm by introducing the Cooperative Network Architecture (CNA), a framework in which object representations emerge from dynamic interactions among recurrently connected neurons. Rather than encoding objects as distinct activation vectors, CNA organizes sensory information into “nets” – structured networks of mutually supporting neurons. These nets are assembled from overlapping “net fragments,” which encode frequently co-occurring feature constellations and are acquired through Hebbian plasticity (Hebb, 1949). This compositional structure supports robust recognition, denoising, and figure completion through cooperative dynamics without external supervision.
In this paper, we present the first concrete implementation of CNA and demonstrate its application to a simple vision task. Our contributions include
  • a formal mathematical definition of nets and net fragments,
  • a computational framework for learning net fragments from sensory input,
  • an illustration of how these fragments can be flexibly composed into object-spanning nets,
  • and an analysis of the emergent filtering and generalization properties of nets.
Our results show that CNA offers a biologically inspired computational framework that integrates local and global information processing. It provides a foundation for more robust and adaptable perception systems, aligning with theoretical proposals for dynamic, network-based neural coding (von der Malsburg, 2018; von der Malsburg et al., 2022). By establishing the neural code and core computational mechanisms within a single representational area, this work provides the essential groundwork for future architectures in which multiple areas cooperate to support invariant object recognition – thus advancing toward the full potential envisioned in network-based neural coding models of perception.
2 The Net Fragment Framework
Before discussing the computational framework, we first introduce its building blocks: neurons, activations, nets, and net fragments. These definitions provide the basis for the discussions in the subsequent sections. For simplicity, we limit the definitions to a single area even though it could be extended to multiple inter-connected areas.
Neurons.
A neuron $i$ is denoted by $n_{i}$. Each neuron $n_{i}$ has a local neighborhood, denoted by $N_{r}(n_{i})$, which is the set of neurons $n_{j}$ such that the distance $d(n_{i},n_{j})$ between $n_{i}$ and $n_{j}$ is less than or equal to a threshold radius $r>0$, i.e.,
$N_{r}\left(n_{i}\right)=\{n_{j}\mid d(n_{i},n_{j})\leq r\}.$ (1)
At each time step $t$, the activity state of a neuron $n_{i}$ is represented by $y_{i,t}\in\{0,1\}$, where $y_{i,t}=1$ indicates that the neuron is active at $t$. The activity of a neuron over time is characterized by an activation pattern, which is a sequence of activation states $\{y_{i,t}\}_{t=1}^{T}$, where $T$ is the total number of time steps (sequential processing steps).
Connections.
Neurons interact via connections. A direct connection from neuron $n_{i}$ to another neuron $n_{j}$ is defined by a positive weight $w_{i,j}>0$, which facilitates the propagation of neuronal activity from $n_{i}$ to $n_{j}$, so that the activity $y_{i,t}$ of neuron $n_{i}$ can influence the activity $y_{j,t+1}$ of neuron $n_{j}$ at the subsequent time step. In this work, we limit the connections to local neighborhoods:
$\forall w_{i,j}\quad n_{j}\in N_{r}(n_{i})$ (2)
Additionally, we limit the weights to a range between 00 and $1$, allowing neuron $n_{i}$ only to positively influence $n_{j}$:
$w_{i,j}\in[0,1]$ (3)
We define a relation $n_{i}\sim n_{j}$ between two neurons to indicate that their activation patterns are highly correlated, i.e., $y_{i,t}$ often influences $y_{j,t+1}$. Under Hebbian learning (Hebb, 1949), the connection strength $w_{i,j}$ increases when there is $n_{i}\sim n_{j}$.
Net Fragments.
A net fragment is defined as the set of neurons that provide direct activation support to a given neuron. In the context of this work, a neuron $n_{j}$ is said to provide support to another neuron $n_{i}$ if two conditions are met: (1) there exists a high correlation between two neurons $n_{j}$ and $n_{i}$ such that $n_{j}\sim n_{i}\Rightarrow w_{ji}>0$; and (2) neuron $n_{j}$ was active at the previous time step, i.e., $y_{j,t-1}=1$. Since synaptic weights are bounded within the interval $w_{j,i}\in[0,1]$, the degree of support a neuron can provide is likewise restricted to this range.
Formally, the net fragment associated with neuron $n_{i}$ at time $t$ is given by
$\mathcal{F}_{t}(n_{i})=\{n_{i}\}\cup\{n_{j}\mid w_{j,i}>0\land y_{j,t-1}=1\}.$ (4)
In this formulation, any neuron $n_{j}\in\mathcal{F}(n_{i})$ contributes to the activation likelihood of $n_{i}$ at time $t$. Notably, neuron $n_{i}$ is part of its own fragment, reflecting a form of self-support – if active at time $t$, the neuron may contribute to sustaining its own activation at the subsequent time step. Note that $n_{i}\sim n_{j}$ is not an equivalence relation since correlation is not transitive, meaning that one neuron can be part of multiple net fragments.
Activation Function.
The decision of whether a neuron becomes active is based on the amount of support it receives from within the net fragment. We call this accumulated input the membrane potential at timestep $t$, which is a function $\mathcal{M}$ of the net fragment $\mathcal{F}_{t-1}(n_{i})$ of a neuron $n_{i}$ at timestep $t-1$.
$a^{\mathcal{M}}_{i,t}=\mathcal{M}\left(\mathcal{F}_{t}(n_{i})\right)$ (5)
The membrane potential $a^{\mathcal{M}}_{i,t}\in\mathbb{R}$ is large when many other neurons provide support, meaning that they “confirm” the plausibility of a neuron’s activity by being active themselves and exhibiting a connection that has been strengthened through previous mutual activity and Hebbian updates.
The activation of a neuron at timestep $t$ is a function of the membrane potential $a_{i,t}$ and environmental conditions $\mathcal{E}$:
$y_{i,t}=\mathcal{A}\left(a_{i,t},\mathcal{E}\right)\in\{0,1\}$ (6)
The environment conditions $\mathcal{E}$ include an attenuation signal, a constraint that selectively suppresses weak excitatory inputs by requiring neurons to accumulate a sufficiently large membrane potential before firing.
Nets.
Neurons are not only influenced by direct connections within net fragments but also by indirect connections (e.g., if $n_{i}$ supports $n_{j}$ and $n_{j}$ supports $n_{k}$, then $n_{i}$ indirectly supports $n_{k}$). A set of directly and indirectly connected neurons is called a net, implementing the basic structure used in this work to represent objects. For simplicity in the following definitions, we introduce $\mathcal{F}^{*}_{t}(n_{i})$, a net fragment $\mathcal{F}_{t}(n_{i})$ in which $n_{i}$ is active:
$\mathcal{F}^{*}_{t}(n_{i})=\mathcal{F}_{t}(n_{i})\quad\text{with}\quad y_{i,t}=1$ (7)
To account for both direct and indirect influences on a neuron’s activation, we use a recursive formulation of nets from the perspective of a single neuron. The base case is defined as the immediate net fragment:
$\mathcal{N}_{t}^{1}[n_{i}]=\mathcal{F}^{*}_{t}(n_{i})$ (8)
For $r\geq 2$, the recursive relation is given by
$\mathcal{N}_{t}^{r}[n_{i}]=\mathcal{N}_{t}^{r-1}[n_{i}]\bigcup_{n_{j}\in\mathcal{N}_{t}^{r-1}[n_{i}]}\mathcal{F}^{*}_{t}(n_{j})$ (9)
The complete network accessible from $n_{i}$ is then defined as the closure
$\mathcal{N}_{t}^{*}[n_{i}]=\bigcup_{r=1}^{\infty}\mathcal{N}_{t}^{r}[n_{i}],$ (10)
which comprises all neurons in the active subgraph that are reachable from $n_{i}$. This recursive formulation illustrates how local net fragments, which can be interpreted as elementary features, are progressively integrated to form higher-order features. Moreover, the explicit time dependence in these definitions underscores the dynamic nature of both the net fragments and the overall network, as they are continuously reassembled in response to the evolving support received from neighboring neurons, and increasing attenuation that prevents neurons with a low membrane potential from firing (see next section).
In the proposed architecture, no symmetry constraint is imposed on the weights, that is, $w_{i,j}\neq w_{j,i}$ in general. However, due to the temporal dynamics of the network, activations tend to stabilize over a longer time horizon $T$, such that for sufficiently large $t$, the activations become approximately stationary, i.e., $y_{i,t}\approx y_{i,t+1}$. Under these conditions, mutual influence between neurons becomes symmetric, meaning that $n_{i}\sim n_{j}$ implies $n_{j}\sim n_{i}$, and almost symmetric weights emerge naturally over time, even in the absence of explicit symmetry constraints.
As a result, it can be assumed that for any pair of distinct neurons $n_{i}\neq n_{j}$, their closures either converge to the same set of neurons or remain largely disjoint. Formally, this can be expressed as:
$\mathcal{N}_{t}^{*}[n_{i}]\approx\mathcal{N}_{t}^{*}[n_{j}]\quad\text{or}\quad\mathcal{N}_{t}^{*}[n_{i}]\cap\mathcal{N}_{t}^{*}[n_{j}]\approx\emptyset.$ (11)
3 Cooperative Network Architecture